# Jianwei Li & Jung-Eun Kim

### v2

@@ B.6 Details of Redundant Units and Alignment Budget
B.7 Parameter-Efficient Fine-Tuning (PEFT) Comparisons
@@
    Prefix Tuning (Li & Liang, 2021): In this approach, a set of continuous task-specific vectors (prefixes) are prepended to the input of each transformer layer, while the rest of the model remains frozen. This method focuses on optimizing only the prefix parameters. In our experiments, we set the learning rate to $10^{-2}$, with a batch size of 16, and fine-tuned for 1 epoch on the Alpaca dataset.
@@
Appendix C Additional Experimental Results
Other Model Families and Downstream Datasets. We include results for Mistral-7B-Instruct-v0.2 and the GSM8K math dataset to analyze the impact of fine-tuning attacks on different model families and datasets. Table 7 presents the pruning results of Mistral-7B-Instruct-v0.2 across safety and utility benchmarks, highlighting the importance of UCU and SCU to utility and safety, respectively. The safety performance of Meta-Llama2-7B-Chat under GSM8K fine-tuning attacks is shown in Table 8, while the results for Mistral-7B-Instruct-v0.2 under Alpaca and GSM8K fine-tuning attacks are reported in Tables 9 and 10. These tables demonstrate that our method is more effective on models that are already strongly aligned, such as LLaMA-2. This finding aligns with previous studies showing that while Mistral-family models outperform LLaMA-2 models on downstream tasks, they exhibit weaker safety alignment and are more susceptible to fine-tuning attacks. Although the improvements of our strategy on Mistral are less pronounced than those on LLaMA-2, it is important to note that Mistral’s initial ASR—measured by LLaMA3-Guard, a relatively more accurate evaluator—is significantly higher, averaging 42.5% across the Adv and HEx-PHI datasets, compared to LLaMA-2’s initial ASR of just 1%. Given this observation, the results are reasonable, as it is unrealistic to expect a model with inherently weaker safety alignment to maintain strong safety performance under fine-tuning attacks. Nevertheless, our method achieves a relative 30% reduction in ASR on Mistral under Alpaca fine-tuning attacks. Notably, our conclusions follow a common-sense premise: the primary objective is to preserve the safety performance of well-aligned models when subjected to fine-tuning attacks.
@@ Appendix C Additional Experimental Results
Additional Jailbreak/Red-Teaming Attacks. To evaluate the impact of finetuning attacks on model robustness against jailbreak and red-teaming attacks, we assess the safety performance of finetuned aligned models under three attack methods: GCG, AutoDAN, and PAIR. Specifically, we utilize the HarmBench framework and sample 120 data points to measure performance. As shown in Tables 11 and 12, our method effectively mitigates performance degradation in adversarial scenarios.
Table 11: Safety performance of Meta-Llama2-7B-Chat under Fine-Tuning attacks (Alpaca) across red-teaming attacks.
@@
Table 12: Safety performance of Meta-Llama2-7B-Chat under Fine-Tuning attacks (Dolly) across red-teaming attacks.
@@
Appendix D Appendix: More Presentations and Discussion
@@
D.1 Ablation Study with Different Ratios of Safety-Critical Components
@@
D.2 Comparing with Previous Work
While previous works have employed similar techniques (Wei et al., 2024b), such as pruning with utility or safety datasets to identify safety-critical components, our approach distinguishes itself in several key aspects:
@@ D.2 Comparing with Previous Work
D.3 Additional Figure Presentations
@@
Figure 13: Conversion of RU (Redundant Units) in MLP module during the alignment process from Llama2-7B to Llama2-7B-Chat across different blocks. The majority of RU remains as RU, with a smaller portion being converted to CU and only minimal conversions into SCU and UCU. This indicates that a large percentage of redundant units remain unused and do not play any meaningful role. Additionally, RU is predominantly located in the middle blocks of the model.
Figure 14: Conversion of $U$ (Utility Units) in MLP module during the alignment process from Llama2-7B to Llama2-7B-Chat across different blocks. This figure illustrates the proportion of $U$ being converted into other roles—CU (Complex Units), SCU (Safety Critical Units), UCU (Utility Critical Units), and RU (Redundant Units)—at various blocks of the model. Most $U$ are converted into CU, with a significant portion also transitioning into UCU and RU. A very small fraction are transferred to SCU. Notably, conversions to CU predominantly occur in the early and later blocks, while conversions to RU and UCU are concentrated in the middle blocks. Although transfers to SCU are minimal, they occur slightly more often in the middle blocks.
Figure 15: Conversion of units into SCU (Safety Critical Units) during the alignment process from Llama2-7B to Llama2-7B-Chat. Most SCU originate from $U$ (Utility Units) rather than RU (Redundant Units), highlighting a reduction in utility-focused units, which contributes to the alignment tax. Meanwhile, a large portion of RU remain inactive throughout the process, representing a significant inefficiency as these units fail to serve any functional role.

### v3

@@ C.6 Details of Redundant Units and Alignment Budget
C.7 Parameter-Efficient Fine-Tuning (PEFT) Comparisons
@@
    Prefix Tuning (Li and Liang, 2021): In this approach, a set of continuous task-specific vectors (prefixes) are prepended to the input of each transformer layer, while the rest of the model remains frozen. This method focuses on optimizing only the prefix parameters. In our experiments, we set the learning rate to $10^{-2}$, with a batch size of 16, and fine-tuned for 1 epoch on the Alpaca dataset.
@@
Appendix D Appendix: Additional Experimental Results
Other Model Families and Downstream Datasets. We include results for Mistral-7B-Instruct-v0.2 and the GSM8K math dataset to analyze the impact of fine-tuning attacks on different model families and datasets. Table 12 presents the pruning results of Mistral-7B-Instruct-v0.2 across safety and utility benchmarks, highlighting the importance of UCU and SCU to utility and safety, respectively. The safety performance of Meta-Llama2-7B-Chat under GSM8K fine-tuning attacks is shown in Table 13, while the results for Mistral-7B-Instruct-v0.2 under Alpaca and GSM8K fine-tuning attacks are reported in Tables 14 and 15. These tables demonstrate that our method is more effective on models that are already strongly aligned, such as LLaMA-2. This finding aligns with previous studies showing that while Mistral-family models outperform LLaMA-2 models on downstream tasks, they exhibit weaker safety alignment and are more susceptible to fine-tuning attacks. Although the improvements of our strategy on Mistral are less pronounced than those on LLaMA-2, it is important to note that Mistral’s initial ASR—measured by LLaMA3-Guard, a relatively more accurate evaluator—is significantly higher, averaging 42.5% across the Adv and HEx-PHI datasets, compared to LLaMA-2’s initial ASR of just 1%. Given this observation, the results are reasonable, as it is unrealistic to expect a model with inherently weaker safety alignment to maintain strong safety performance under fine-tuning attacks. Nevertheless, our method achieves a relative 30% reduction in ASR on Mistral under Alpaca fine-tuning attacks. Notably, our conclusions follow a common-sense premise: the primary objective is to preserve the safety performance of well-aligned models when subjected to fine-tuning attacks.
@@ Appendix D Appendix: Additional Experimental Results
Additional Jailbreak/Red-Teaming Attacks. To evaluate the impact of finetuning attacks on model robustness against jailbreak and red-teaming attacks, we assess the safety performance of finetuned aligned models under three attack methods: GCG, AutoDAN, and PAIR. Specifically, we utilize the HarmBench framework and sample 120 data points to measure performance. As shown in Tables 16 and 17, our method effectively mitigates performance degradation in adversarial scenarios.
Table 16: Safety performance of Llama2-7B-Chat under Fine-Tuning attacks (Alpaca).
@@
Table 17: Safety performance of Llama2-7B-Chat under Fine-Tuning attacks (Dolly).
@@
Appendix E Appendix: More Presentations and Discussion
@@
E.1 Ablation Study with Different Ratios of Safety-Critical Components
@@
E.2 Comparing with Previous Work
While previous works have employed similar techniques (Wei et al., 2024), such as pruning with utility or safety datasets to identify safety-critical components, our approach distinguishes itself in several key aspects:
@@ E.2 Comparing with Previous Work
E.3 Additional Figure Presentations